OpenStudy (anonymous):

limit x-> 5 sin(x^2-25) / (x-5)

OpenStudy (lgbasallote):

that's equal to lim x => 5 sin(x-5)/(x-5) times x+ 5 sin(x-5)/(x-5) = 1 according to many theorems so.. 1(x+5) = x+5 substitute 5.. 5+5 =10

OpenStudy (anonymous):

10

OpenStudy (anonymous):

thanks alot.

OpenStudy (anonymous):

the above solution of lgbasa. is wrong

OpenStudy (anonymous):

because you can't write sin(a*b)=(sina)*b

OpenStudy (anonymous):

so, what's the correct answer ?

OpenStudy (lgbasallote):

it's sin[(x+5)(x-5)]/x-5..you can't apply double angle theorems or whatever theorems there so i think my solution is valid

OpenStudy (anonymous):

\[\sin(x^2-25)/(x-5)=(x+5)*\sin(x^2-25)/((x-5)(x+5))=(x+5)*\sin(x^2-25)/(x^2-25)\]\[x->5 => x^2->25\]\[\lim_{x^2 \rightarrow 25}\sin(x^2-25)/(x^2-25)=1\]so (5+5)*1=10

OpenStudy (anonymous):

check this out

OpenStudy (anonymous):

that's the correct answer dipankarstudy ?

OpenStudy (anonymous):

yes i think so...you can consult others..but definitely you can't do the first one

OpenStudy (anonymous):

ok, thnx